OpenStudy (anonymous):

If the sum of the measures of the interior angles of a convex polygon is 5400, how many sides does the polygon have?

OpenStudy (jdoe0001):

sum of interior angles for a regular polygon = \(\bf (n - 2)\times 180\) thus, since we know that that sum is 5400, then \(\bf 5400 = (n-2)\times 180\) n = number of sides solve for "n"